5.​ Obtain the value of P3 (x) using Rodrigues' formula​​​​​       

Pn(x) = 1/(2n)n! x dn/dxn *(x2-1)n

1
Expert's answer
2022-04-27T13:34:52-0400
"P_3(x) = \frac{1}{(2^3)3!} \frac{d^3}{dx^3} (x^2-1)^3"

"= \frac{1}{48} \frac{d^2}{dx^2} (6x(x^2-1)^2)"

"= \frac{1}{8} \frac{d}{dx} ((x^2-1)^2+4x^2(x^2-1))"

"= \frac{1}{8} (4x(x^2-1)+8x(x^2-1)+8x^3)"

"= \frac{1}{2}(5x^3-3x)"